January 24, 1996
CCD Imaging Spectrometer
George C. Marshall Space Flight Center
National Aeronautics and Space Administration
Marshall Space Flight Center, AL 35812
Center for Space Research
Massachusetts Institute of Technology
Cambridge, MA 02139
This report covers the period November 1995.
The monthly status meeting for ACIS was formally conducted on November 30, in conjunction with the ACIS mid-term review which was conducted on November 29. The meeting was attended by Max Rosenthal from the MSFC Project Office and Harvey Tanenbaum from SAO. Due to the mid-term review the previous day, Charlie Jones (AXAF project engineer) and several MSFC technical personnel were also in attendance. In addition to the usual topics of MIT schedule status (Mayer), MIT technical status (Goeke), and the science activities (including the latest CCD status) by Mark Bautz, the presence of the LMA personnel at MIT for the mid-term review allowed a rather thorough review of the PSMC technical and schedules status, presented by Lloyd Oldham, Ed Sedivy and Brent Reschak. The next monthly review is scheduled for the week of January 15 at NE80. The exact date has not yet been set.
The monthly Technical Interface Meeting with LMA was conducted at MIT on November 28 in conjunction with the mid-term review and monthly status review the following days.
MIT participated in the AXAF project level telecons on October 31, November 7, 21, and 28. There was no AXAF telecon on November 14 due to the shutdown of the Federal Government for this week.
MIT participated in the ACIS bi-weekly status review on November 1. The telecon normally scheduled for November 15 was also canceled due to the government furlough, and status review for November 29 was conducted in person (ACIS mid-term review).
During the month of November, the overall ACIS schedule was continually updated and expanded. The status was presented at the meeting on November 30. At that point in time approximately 120 tasks had been completed; approximately seven tasks had been completed behind the schedule established on August 1. Due to the fixed nature of the schedule, many of the late tasks are the result of the lateness of the PCB layout of the DEA Analog Board. The activities for the loading and testing of the first calibration PCBs were re-programmed to recover some of the lost time in the critical path. At the end of November, delivery is still scheduled for December 14, 1996, with a total of seven weeks of contingency remaining. The option of performing the DEA/DPA/Detector Assembly thermal vacuum testing at Lincoln Lab as part of the existing System Level test is under serious consideration in an effort to shorten the critical path and increase the slack on this delivery date. The down side of this option is that the flight PSMC would not be available to participate in this combined test program. MSFC is considering the effects of this option and an Integration and Test meeting has been requested by the MSFC project office in the Jan.-Feb. time frame.
During the monthly TIM at MIT on November 28, LMA did report that they had worked out the details of a procedure for the alignment of the Detector Assembly, which fit into the available time slot identified during last month's TIM. Therefore, the alignment activity will not affect the overall delivery date of the ACIS experiment.
The ACIS project remains in relatively good shape from a schedule point of view. In fact, enough tasks have now been completed that resource deficiencies (thermal, power from the PSMC, and weight) have taken over as the driving problems with the ACIS experiment. One of the main events in November was the MSFC request at the beginning of the month for MIT to perform a complete review of the ACIS weight status. A report was submitted on November 6 which showed an ACIS `current' weight of 296 lbs (vs the 260 lbs allocated). This obviously caused quite a stir at MSFC. Max Rosenthal sent E-mail to the ACIS and HETG science team members requesting a science reduction plan to get the ACIS weight back to 260 lbs. Internal ACIS/HETG meetings on this topic were conducted on Nov. 16 and 21. In the meantime, the MSFC project office formed an ACIS Mass Audit Review team, led by Jack Loose of MSFC, to try to determine the validity of the ACIS mass estimates provided by MIT. Background data was immediately provided to the review team and a mass audit was scheduled as part of the ACIS mid-term review on November 29 (and carried over to November 30). When all was said and done, the best estimate for the ACIS weight was 293 lbs.
During November, no personnel changes were made to the ACIS staff at either MIT or LMA. On a positive note, Dorothy Gordon, who had earlier indicated that she was seriously considering leaving MIT, has decided to stay on as the DPA digital electrical engineer for the foreseeable future.
Seven devices were received in November; six of these were flight candidates. The seventh device, though determined by Lincoln before shipment to have been damaged during packaging at Lincoln, was delivered as a facility test device.
All five of the front-illuminated flight candidates delivered during the month passed the screening evaluation test. The single back-illuminated flight candidate delivered during November failed to pass the screening test on account of its high dark current.
Flight CCDs, which have passed screening tests as of November 30, 1995 and are available for use in the flight Focal Plane, are listed below:
|Device||(FI/BI)||Date Rec'd||Date Screened
|w147c3||BI||10 Aug.||11 Aug.|
|w129c2||FI||15 Sep.||15 Sep.||Shows "D noise"|
|w87c4||FI||3 Nov.||3 Nov.|
|w78c1||FI||7 Nov.||8 Nov.|
|w97c1||FI||17 Nov.||17 Nov.||Shows "D noise"|
|w157c1||FI||24 Nov.||24 Nov.|
|w157c2||FI||24 Nov.||27 Nov.||Shows "D noise"|
Performance data for these devices is available on the ACIS web site. It was determined that the D noise problem is not confined to any particular lot of flex prints. Kissel discovered that the D noise can be reduced to low levels (4.3 electrons, RMS, or lower) by tuning the output transistor drain bias.
In order to increase the flux available from the tritium calibration source, a significant modification of the quantum efficiency chambers was undertaken. The modification will allow the source to be placed closer to the devices under test.
Revisions were made to increase the stability of the temperature control in the HIREFS chambers.
The radioactive source for the ACIS contamination monitor (internal, door-mounted source) was specified. The source will be Fe55 of low activity (10-50 micro-curie), with a thin protective coating (nominally 200 micro-gm/cm^2 nickel) to allow detection of both L and K characteristic X-rays.
|8 Nov.||Calibration Implementation Team, Huntsville|
|28 Nov.||TIM with LMA, Cambridge|
|30 Nov.||2-chip ACIS discussions with MSFC, SAO, in Cambridge|
The ACIS Contamination Monitor (Door Source) design and details are complete. An engineering unit has been built. Some refinement is required, including weight reduction to reduce the loads on the Door.
The External Calibration Sources are being defined by the Scientists.
The vendor is working, but delivery has been delayed.
No activity at this time.
Design continues. Details of the structure are nearly complete. The DEA to DPA cable routing has been addressed. ACIS has not been able to contribute to solving the assembly problem without changes which will affect the SIM interface. The design effort is working with the existing interface.
No activity during this reporting period.
No change during this reporting period.
Acoustic and vibration testing has been successfully completed at LMA on the thinnest samples.
The lifting fixture for the Support Structure, DEA, and DPA has been redesigned and the details are ready for release.
Detail design of the DPA engineering unit is complete. Detail drawings of the housing have been reviewed and are in final engineering checking and update. Purchase orders have been placed with two outside machine shops for fabrication of the dip brazed -Y panel, the other panels, PCB frames and covers. The CSR machine shop will fabricate the small brackets and other small pieces necessary to complete the assembly. The DPA circuit boards, Front End Processor, Back End Processor, processor preparatory, Backplane and power/sync distribution are in the review processes.
Completed the Mid-Term Review and weight audit. Presently evaluating and responding to the action items generated from the meetings.
Refinement of the ACIS cable interconnection continues. Particular attention has been given to the DPA/DEA/Support Structure housekeeping lines.
Supported 11/2 telecon on thermal requirements for BASD, for the ITCS (ISIM Thermal Control System).
Participated in Thermal TOP telecon on 11/30.
Participated in thermal discussions at LMAC on 11/7 concerning PSMC redesign options.
The ACIS thermal models were revised to decrease the DPA hot-case temperatures with or without an opening in the SIM wall, and to reduce heater power required in cold cases. The -Y wall of the DPA was changed from black to white paint. Coatings on the +Y and -X walls of the DEA were changed from black paint to irradiate.
Drafted the ACIS Thermal Vacuum Test Plan and discussed it with MIT and LMAC engineers during the MIT/LMAC TIM on 11/28.
Participated in the ACIS Mid-Term Review at MIT on 11/29-30, including the Thermal splinter meeting. Delivered to TRW and MSFC thermal engineers was the latest ACIS thermal modeling information to be incorporated in BASD thermal models.
The first engineering board was fabricated this month. Component population of the board was also completed. Engineering activity this month focused on the debug and testing of the first engineering board.
The BEP (Back End Processor) and its interfaces are being debugged using the flight test software. Some minor bugs were found in the command interface and fixed. Support of the software group will continue until the BEP hardware passes all flight software diagnostics.
Our consultant, Jim Littlefield, implemented test software which allows taking GSE script files from the DEA and sending them through the BEP to the DEA interface.
The interfaces to the FEP and flight ROM's are being debugged.
Back End Processor schematics have been updated and are in the review cycle for next release.
The Front End Processor: Thresholding and parity plane checking have been turned on in conjunction with memory bus "banging", effectively testing most of the hardware subsystems in simultaneous operation. Most of the diagnostics are complete and hardware checkout continues.
The DPA Backplane is ready for operation. We expect to integrate the Front End and Back End Processor in about a week or two. The ICD for the BEP/FEP Interface was approved and released to Rev. A.
The PREP and Power I/F board schematics have both been completed and are awaiting minor revisions for release to Rev. A. There have been some systems changes, which are now beginning to solidify, involving interconnect and power distribution. These changes (which also impact the Backplane, FEP and BEP boards) will be incorporated in the PREP and Power I/F boards.
Frame Buffer circuit board #1 (configured to generate continuous ramp frames) has not been updated as yet and is still operating with a FEP. Testing of Board #2 has been completed and the board has been installed in a temporary enclosure. Board #3 is awaiting testing. The panels for all three enclosures have been machined and will be assembled and wired as the circuit boards become available.
Reviewed and commented on the XC05 XRCF Interface Definition for EGSE issues. Provided software and trouble shooting assistance for Frame Buffer testing and development. Worked with Bldg. 37 staff to develop a common FITS header file that can be used for image capture as part of CCD and CCD controller testing. Continued to work with the ACIS flight software group to develop test tools.
Published Rev. C of the Software Requirements Specification, and prepared a copy without change bars for MSFC baseline review.
Closed 14 of 18 outstanding tracked comments from CDR. Of the remainder, three await responses to our proposed implementation and the fourth is a question of documentation which will be addressed in the next delivery of the ACIS S/W Detailed Design Specification in March 1996. They are all visible on the ACIS Web site: "http://acis.mit.edu/swmtr/".
We asked the review team (Crumbley, Purinton & Tennant of MSFC, Yester of TRW) to assist us in working with Rogson (TRW) to finalize the IP&CL binary field specification so that we can submit a definitive description of ACIS command, parameter block, and telemetry packet fields.
Conducted ECO review of Science Manager and Data classes.
Tested BEP serial command interface; several problems detected require changes to BEP firmware. Began testing BEP telemetry interface. Received updated specifications of DMA Image Loader. Continuing to test high-level software in FEP.
A pre-alpha-release was given to ACIS Q/A on November 27.
All flight software is now being subjected to unit and coverage tests.
Attended meetings of the Test Group, which reviewed an updated version of the Test Tool Specification written by MIT.
Six (6) Alerts from NASA/MSFC were received over the report period. These items are listed below. They were compared with the available MIT parts lists, none of the Alerts impact the MIT ACIS design at this time.
|ALERT #||MSFC #||Part Number||Part or Material Name|
|NONE||6850||ALL SSDI PARTS||ALL MIL-S-19500 DIODES|
|FJ-P-96-01||6886||MS27130-CR8K||NUT, BLIND RIVET|
Revised Bill of Material (BOM) for DEA, produced documentation and labels for three (3) DEA engineering unit kits.
100% PIND, 100% X-ray, and sample DPA for flight microcircuits, transistors, and diodes continues at Associated Testing Laboratories, Inc. (ATL), in Burlington, MA. The total number of flight part types screened to date is forty-eight (48).
Preseal visual examination was done at Space Electronics, Incorporated (SEI) on ninety-one (91) pieces of the (ninety-one (91) Field Programmable anti-fuse (8000) Gate Arrays (FPGAs), P/N A1280ARPZY.
Seven (7) items were sent to MSFC for 1443 testing. These items are:
Coordinated the cleaning of five (5)1" OWS holders at MIT Lincoln Labs.
Eight (8) OWSs, which were received from MSFC, were measured for baseline NVR thickness.
Lockheed Martin Astronautics (LMA) submitted five (5) waivers against the requirements of NHB5300.4 (3A-1), NHB5300.4 (3J), NHB5300.4 (3G), NHB5300.4 (3H), NHB5300.4 (3I), and NHB5300.4 (3K). The reasons stated by LMA for these requests is that they would like to use the LMA procedures rather than the NASA procedures. In the case of NHB5300.4 (3I), LMA stated that they could not find a manufacturer willing to build in accordance with the NHB requirements and the cost would be prohibitive. MIT provided the names of three (3) manufacturers willing to build to the NASA requirements and rejected the waiver request. MIT submitted the other LMA requests as deviation requests 36-002, 36-003, 36-004, and 36-006 for MSFC consideration.
MIT has been discussing design and fabrication of printed circuit boards (PCBs) in accordance with the requirements of the NASA Handbooks (NHBs). The deviation granted to TRW in this regard has been incorporated into MIT specification 36-02105, revision D. This document has been sent to potential PCB suppliers to determine if we have a reliable source for flight boards.
SCD 36-02311 was prepared and released for a radiation hard 38.4 MHz crystal oscillator. A NSPAR on this part was prepared and submitted to MSFC for approval. A purchase order was placed with Q-Tech for four (4) flight crystal oscillators and three (3) engineering (commercial) units.
Prepared SCD 36-02350 for an Fe55 source. RFQs were sent to Amersham, Isotope Products Laboratories (IPL), and DuPont Merck. DuPont Merck submitted a "no-bid," but a quote has been received from IPL. Amersham will submit a quote by 12/8/95.
One (1) new NSPAR was submitted for approval on the crystal oscillator. Approvals were received on six (6) previous submittals.
|36-002||A to D Converter
|36-003||CA Memory Module
|36-004||FB Memory Module
|36-005||Programmable Supply current
Op Amp 36-02304
Only Memory 36-02306
|36-008||Electrical Connectors, PCB
Mount SND Type
|36-009||Electrical Connectors, PCB
Mount KA Type
|36-012||Junction Field Effect Transistor
|36-013||Dual Surface Mount Diode
|36-014||Dual Operational Amplifier
|36-015||8000 Gate Anti-fuse Field
Programmable Gate Array
|36-017A||Charge Coupled Device
|36-018||Microcircuit, Octal Buffer
|36-019||Microcircuit, Octal Bus
Transceiver (Harris HCS245)
Flip-Flop (Harris HCS374)
Line Driver (Harris HS26C31)
Differential Line Receiver
Q-Tech part type
Lockheed Martin Astronautics (LMA) submitted twenty-six (26) NSPARs to MIT. These were reviewed and all sent back to LMA with comments.
Radiation testing has been completed at Space Electronics, Inc. (SEI) on thirteen (13) device types. Results of these tests are listed below.
|MFR.||PART NUMBER||RADIATION TEST
|Analog Devices||DAC8800BR/883||<2K Rads|
|Com Linear||CLC505A8D||>100K Rads|
|Harris (Chip Supply)||36-02305 (CA 3080)||<100K Rads|
|Analog Devices||OP220AJ/883 (TO-5 can)||8K Rads|
|Texas Instruments||TL082/883B||>100K Rads|
|Analog Devices||REF43BZ/883||>200 K Rads|
|Chip Supply||OP-220 (DIP)
Devices which have not passed 100K Rads of Cobalt 60 testing will be shielded or design work-arounds will be implemented. Eight (8) more device types are planned for radiation testing. These are listed above without results.
Letters of interpretation updated.
Updated JSC 44 Form for back Door and sent it to H. Hooper.
Reviewed Hazard Report Package from TRW, to B. Mayer. All of the issues have to do with GSE operation at KSC. Extracted applicable information and sent it to B. Bond of LMA.
There has been no activity on the Performance Assurance and Safety (PAS) Plan. The PAS Plan in effect is revision B.
A complete vacuuming and cleaning was performed in the CCD assembly tent to prepare for flight testing. Now that flight testing has started, the washable ESD cleanroom suits are being worn.
Started pursuing possible candidates for cleaning and 1238 certifying the X-ray source.
Drafted a new procedure for installing the CCDs and optical witness samples (OWS). This is still under review.
An ECO was generated for the shipping and handling procedure for CCDs.
Drafted procedures for bagging material.
Drafted procedure for measuring silicon witness samples for contamination.
In process of developing a new monitoring procedure for CCDs during vacuum testing.
Received the RCL meter from calibration at Draper Labs.
Five vacuum chambers at building 37 for ACIS CCD calibration are waiting to be certified using the optical witness samples.
Performed NVR and particle measurements on 6 locations being monitored. Monitoring of particle levels in clean tents showed an unacceptable increase in particle fallout in the Hetterick clean tent for a four week period. Normal fallout rates range from 0 to 100 particles for periods of two to four weeks. The Hetterick clean tent had a fallout rate of 2,649 particles. Investigation proved the root cause to be from major troubleshooting by the scientist to find a noise problem. The tent was shut off during parts of the troubleshooting. The tent has been re-cleaned. No flight CCDs were exposed to this condition. Some excessive particulate generation was also inside the CCD assembly clean tent in building 37, due to installation of shades in the windows. The tent has been cleaned and no flight CCDs were exposed to this condition. The assembly areas are seeing an NVR buildup of about 4-8 Angstroms per month.
The two inch OWS designed for monitoring contamination on CCDs can only be used 50 % of the time due to interference if two CCDs are tested simultaneously. We have developed 1" OWS holders that will hold a 1" MgFl OWS to the framestore cover, to each CCD that will allow the monitoring of contamination on each CCD.
The large vacuum chamber at Lincoln Laboratory is being analyzed for cleanliness to perform three tests for ACIS; Thermal Cycle, Science Test, and 1238 Certification. A TQCM, along with two 4" silicon witness samples were installed to obtain data for particulate and NVR contamination.
The NVR level of contamination in the Lincoln Lab chamber is acceptable for the flight science test and thermal cycle testing. The particulate level of contamination is unacceptable at this time, however, We believe that the chamber can be cleaned to meet our criteria prior to the flight tests. 1238 certification analysis has not been done yet.
Transmission tests on engineering OBFs were performed at the Brookhaven National Laboratory (BNL) in Upton, New York on Nov. 19 and 20. OBF platens were received by MIT from Dr. George Chartas of Penn State, who performed the transmission tests. A fit check proved some redesign needed to be done. After they were reworked, the platens were cleaned by Lincoln Laboratory and sent to BNL. MIT Contamination Control Engineering supported the installation of the OBFs into their beamline vacuum chamber, and analyzed the chamber for cleanliness. The installation process was risky. Some of the problems were: 1) The OBFs were exposed for an excessive amount of time throughout the installation process; 2) Grease is used on bolts that hold the conflat flange to the chamber. These bolts are removed using the gloves in the glove box which will be the same gloves used to install the OBF fixture into the chamber. The risk of cross contaminating the fixture and the frame is high. It may be possible to place another glove over these glove boxes, but this would further desensitize "feel"; and 3) The actual installation of the fixture into the chamber is very difficult. The size of the OBF fixture is larger than the opening in the feed through box and the chamber, so twisting and manipulating the fixture is necessary. Results for contamination tests at BNL showed an unacceptable increase in NVR. The witness samples used had an increase of 12 Angstroms over a 36 hour period.
Transmission tests on engineering OBFs were performed at the Wisconsin Synchrotron, in October. Some open issues remain. The spectrometer OBF fixture has been redesigned and sent to PSU to rework for the next scheduled trip to Wisconsin. The Imager OBF fixture design is in process. The results from the witness samples and contamination monitoring reveal that this chamber is clean for future flight filters so long as a few improvements are made. A class 100 clean tent installed around the chamber, along with adding a vacuum pre-bake prior to filter installation, will ensure that the process will pose minimal contamination to the OBFs.
Participated in review of 3 software modules: Science Data Processing Class, Science Management Class, Start-up Class.
Participated in the review of the EGSE test tools being developed to test and verify the ACIS Instrument.
Began execution of test scripts to verify memory read, write and execute capabilities. Test being performed on simulated hardware.
Have received and built the integration version of the flight software.
Participated in technical meeting with MSFC. Agreed to supply them with a sample of test procedures and a sample traceability matrix by the end of 1995.
During this month, five flight candidate FI chips, one non-flight FI chip, and one flight candidate BI chip were shipped to campus, resulting in a total of seven devices delivered this month, in excess of the goal of six parts. All five of the FI flight candidates were considered acceptable, raising the total number of FI candidates accepted by CSR to six. The non-flight part was included for tests of output D noise problems, and the results have yet to be obtained. One hypothesis is that the output D noise has its origins in the flex print. We have shipped parts that have flex prints made in batches 3 and 4 from Flex Tech and 1/2 of the batch 3 parts showed noise, while 0/2 parts from batch 4 showed noise. These statistics are not large enough to draw any conclusions. We are currently pursuing the possibility that the noise may be associated with a stub on the D output line that is required for gold plating of the flex prints. This stub has been manually removed from several flex prints as a test.
The BI chip was rejected due to high CTI after testing at CSR to -110C. This chip was from BI Lot 4, which was found to have high dark current at -50C during final testing at Lincoln. This result suggests that the two other potential flight chips from this lot, which have similar Lincoln test results, will also have high CTI. Therefore, not counting these devices, we have eight more unpackaged BI chips from lots 1 and 2 as flight candidates, or a total projected quantity of four packaged BI chips that could be incorporated in flight paddles.
We remain on schedule for delivery of parts to campus and will attempt to deliver six to eight chips during December. Presently, these are scheduled to be FI parts, since BI part packaging is on hold until a decision is reached by CSR on what flex prints to use for these devices because of the D noise question.
Parts other than chips, that is, flex prints, tees, and alumina blocks, are in adequate supply for the present. The next, final batch of flex prints (70) is due from the vendor early in January.
Metallization of the flex print circuits for the low thermal emissivity coatings continues to proceed smoothly. The vast majority of the available flex print circuits have been metallized, all of them successfully, without the failures which plagued us this fall. The remainder will be completed during the first half of December.
Environmental testing of the Optical Blocking Filters (OBFs) in the CAMSIM Camera Assembly has been completed at LMA. The Detector Assembly is expected to be returned to Lincoln at the end of January for evaluation. All indications are that the CAMSIM Detector Assembly survived this additional testing.
Eddy current testing is performed on the flight beryllium structures to ensure that no cracks are present. A custom probe for the eddy current testing of the spotface repair was fabricated and successfully tested on the master defect block. The flight beryllium structures are scheduled to be tested during the first week of December.
Five material samples were submitted to MSFC for 1443 testing. These include Minco heater elements fabricated to the S-322-P-079 specification, high reliability RTDs, marking ink, and metallized Kapton tape.
During this month an engineering estimate for the tasks associated with extending the ACIS delivery date to December 1996 was provided to MIT. This estimate was reviewed by MIT/CSR and comments and task reductions coordinated with LMA. The change order associated with performing the work and preparing a firm estimate was planned to be available in early December with submittal of the firm proposal before Christmas.
The ACIS Program Technical Interchange Meeting (TIM) was conducted at MIT/CSR. The meeting provided detailed technical coordination on the LMA major hardware status and provided a forum to review the PSMC status presentation for the ACIS monthly management status review. The TIM was followed the next day by an ACIS Mid-Term review which provided detailed updates of the ACIS status and technical product updates since CDR. Since TRW was in attendance, this review also enabled discussion of concerns with the AXAF grounding to be initiated. These discussions were not completed within the time available and were planned to be continued at a later date. Concurrent with the second day of the Mid-Term review, the ACIS Monthly Management Status was conducted. LMA supported this review by presenting a detailed discussion on the PSMC schedule delays, the causes of the delays and the lessons learned. This review demonstrated that the PSMC design meets the requirements as demonstrated in engineering unit hardware.
The checkout difficulties with the RCTU approximator, PSMC test tooling, and the PSMC engineering unit (EU) were resolved this month and the EU was ready for EMI/EMC tests. This test was waiting for the EMI/EMC test facility to become available.
Installation of a residual gas analyzer in the 3' x 5' chamber was completed this month. However, the 3' x 5' chamber bakeout and operational checkout was slowed by a leak that developed in one of the chamber rework welds during the extended high temperature bakeout of the chamber. Chamber certification is now planned for December.
PSMC closed frame test configuration was completed this month. Two major system tests remain; thermal cycle and EMI. It is anticipated that the EMI test will be performed the beginning of December and the thermal cycle test will follow near the end of the month. The suite of EMI tests will be structured such that the most important tests are performed first. This will allow maximum time to be spent on the primary issues of concern. Conducted emissions and conducted susceptibility will be performed first, followed by radiated magnetic emissions tests.
As a note, the test problems encountered earlier, as reported in previous status reports, have been cleared up and testing has progressed in a much more predictable manner during the latter part of the month. It is anticipated that the complications of the testing phase (i.e., the RCTU approximator and the DEA/DPA test tool) are primarily behind us.
LMA attended a TIM at MIT on Tuesday, November 28, 1995. The following day various members from MSFC joined us and conducted a detailed PSMC review. Max Rosenthal and Dick Acquire of MSFC and a representative from TRW reviewed the technical details of the PSMC at length. Photographs of the PSMC (current as of the prior week ) were provided and discussed, along with a status of the various portions of EU#1 and EU#2. The most recent set of lab results from the power supplies, such as phase/gain stability plots, regulation data, etc., were also reviewed. This information alleviated any fears concerning the integrity of the PSMC design. On Thursday, LMA staff presented a status review to MIT and MSFC of the PSMC detailing the current problems we have encountered and the lessons learned. Anticipated changes regarding DEA and DPA Power Supply requirement changes were discussed briefly.
Assembly of the second engineering unit (EU#2) continues. This unit will serve as a test bed for LMA after EU#1 is delivered to MIT. Substantial progress has been made in the assembly of the PWBs for EU#2. The Detector Housing (DH) Thermal Controller is 98% complete, Digital Power Supply is 100% complete with initial testing underway, Analog Power Supply is 95% complete, the Serial Digital Telemetry is 100% complete, Vent Valve and Mechanism Controller is 95% complete, the IO/EMI cavity is 100% complete, and the Motherboard is 95% complete. The build of the second EU cable harnesses was started and completed in November. Some changes from the EU#1 harness have been identified and incorporated into the EU#2 cable harness.During the TIM of late November, 20 NSPARs were informally delivered and reviewed by MIT. The comments are being considered.
A detailed radiation shielding analysis was performed which provided a dose map of the inside of the PSMC. Nine points on each printed wiring board (PWB) have been identified and the total radiation levels have been calculated. Individual part types are now being reviewed and analyzed for radiation tolerance. Given that the radiation levels are known within the PSMC and that the radiation tolerance levels are known and understood for a part type, a radiation tolerant PWB layout can now be achieved. For the cases where simple placement is not adequate, additional local shielding will be added. This information will be used in the layout of the flight PWBs.
The Detector Housing details are in the last stages of fabrication. Inserts are being installed and parts are being precision cleaned. A concept for the survival heater connector bracket is available. This new requirement has resulted in a delay to the release of the Detector Housing Assembly drawing which has pushed the planned start of assembly out into January 1996. Resolution of the paraffin leak that occurred during the StarSys actuator life cycle test has also delayed the schedule.
Fabrication of TCS components is on schedule with no major problems to date. The TCS installation drawing will be released in December.
The rest of the development OBFs were received from Luxel. One set of 2000 Angstrom Lexan filters and 2 sets of 3000 Angstrom Lexan filters were received. The OBFs were thermal cycled five times between -76 deg C and +46 deg C. No degradation (visually) was noted between pre and post test.
An OBF integrity monitor was developed and tested for use in the upcoming acoustic and vibrations tests. It included a photo-diode and bracket mounted on the Focal Plane, in place of one of the CCDs. A red LED and bracket have been mounted on the opposite side of the filters, in place of the flight LED/RTD bracket. An amplifier circuit picks up any failure in the filters as the output of the photo diode increases when light passes through holes in the filter. A 0.3" diameter hole in one of the filters is discernible with the amplifier circuit.
A pump down procedure for the Detector Assembly was written and validated. Using the procedure, the Detector was evacuated at less than 12 torr/minute to 21 torr internal pressure. The first three minutes, the Qual-level acoustic test was completed. The monitoring equipment did not show a failure of the 2000 Angstrom OBFs. This was verified with a post test inspection. Vibration testing of the filters will occur in December.
The stress analysis CDRL update is in progress and completion may have to be delayed from the current 1/15/96 promise due to the increased test activity being implemented. OBF test activities and design and drawing maintenance tasks on the TCS drawings have delayed the completion of the stress analysis. There are no stress concerns to note at this time and all components analyzed to date are showing significant positive margins.
Preliminary analyses were performed on the different options related to PSMC design changes in preparation for the MIT/LMA TIM. These analyses showed several areas that need to be evaluated in more detail that may lead to future concerns. These areas will be evaluated in more detail once authorized to proceed is received.
Failure analysis of the failed actuator, which was reported in the previous month's progress report, has been started. The failure has been attributed to an improperly seated o-ring. This allowed paraffin past the low pressure seal and precluded proper function of the high pressure seal, resulting in an eventual leak in the high pressure seal. The failure to properly seat the o-rings resulted from several steps which were not performed correctly as the heater sub-assembly was installed in the actuator body. Although the flight actuators were suspect to having this same problem since they were assembled at the same time; however, X-rays of the actuators showed the o-rings to be properly seated. StarSys data shows that if the o-rings are fully seated, which was verified with Xrays, there will be no problems with the o-rings not sealing. The failed life cycle actuator was re-assembled using the same mechanical components with new o-rings. Using the correct seating procedure, the heater sub-assembly was reinstalled. Proper seating was verified via X-ray and a non-destructive disassembly after the actuator was charged with paraffin. Twenty-five ambient cycles were applied using the life cycle load simulator with no degradation of performance. The actuator was disassembled again, verifying that the o-rings had not moved or leaked. Life cycle testing was restarted in the T-vac chamber at -35íC. Approximately 60 of the 128 cold cycles were completed as of this writing without any degradation in performance as was measured previously when the actuator was not sealing properly. The rest of the 253 life cycles will be completed in December. The final failure analysis report will also be completed and released to MIT/MSFC.
The test reports were completed on the mechanism life cycle tests. The reports will be released after all the StarSys anomalies are resolved and the individual actuator life cycle test is completed.
The venting subsystem tubes have been inspected with dye penetrant. All of the critical welds passed the inspection. Several small flaws were detected on the bracket that is welded to the tube to supply support at the mid point of the long section of the tube. Stress engineering inspected the defects noting the defects to be smaller than the critical length of 0.15" with slight porosity. The condition does not affect the strength of the weld and was dispositioned as acceptable for flight since this weld is structural only and have no vacuum seal requirements. The tubes were sent on for X-ray inspection.
A discrepancy report will be processed after all inspections are complete. Preliminary findings are that the tube can be used as is.
Preparation of the life cycle test procedure for the small vent valves and the test plans for leak checking the small valves, the large valve and the vent tubes were initiated this month. These plans will be submitted for approval in December.
The drawings for the GSE were essentially complete this month. This supports planned release of 90% of the GSE drawings by January 1996.
Testing is continuing on the VGSE #1. The first turbo pump has been received and will be tested in the first set of GSE, in December.
Firmware for 90% of the VGSE subroutines have been completed to date. Testing of these subroutines will be performed in December. We expect that all subroutines will be complete by the end of next month.
The shipping container requirements have solidified. The SIM simulator will now require a smaller container since it will contain the Detector Assembly, Venting Subsystem, DEA, DPA and Support Structure. However, this appears to be larger than any of the containers that are currently available and a new special container will probably have to be designed and fabricated.
The light shade analysis was completed this month. The analysis showed the current design is adequate for all conditions except direct illumination of the aperture by the Earth. This is considered to be an unlikely operating condition. However, John Sharp of MSFC is looking into the likelihood of this event. If Earthshine poses a significant problem during the active part of the orbit we have investigated options to modify the design with minimal impact. The most likely scenario is to paint the large vent valve and the short vent tube black using chemglaze Z306. This increases the attenuation of the venting subsystem by more than a factor of 100. The analysis shows that this will reduce Earthshine to acceptable levels. The paint is a polyurethane cured by baking at high temperature. The paint has been MSFC 1443 tested and passes.
Review of the Contamination Control Plan (CDR Version) is in process at LMA.
A leak in the 3'x5' chamber was discovered late in this reporting period. The source of this leak was determined to be a cracked weld in a flange. Repairs are being planned. Following repair, testing with the newly installed RGA will continue.
The FMEA and CIL update for ACIS was reviewed with MSFC and MIT personnel on 11/27/95. PTS items have been identified and a first cut at the DPS items is in process. A plan to derive DPS inputs was proposed by MIT to have LMA get the needed information from MIT over the next month or so, with incorporation into each document by mid-February and publication by the end of that month. An agreement was reached in the joint MSFC/MIT/LMA meeting to include any software detection and mitigation capability into the FMEA because there is no requirement for a separate "software FMEA" applicable to ACIS.
The EMI/EMC test procedure for tests of the engineering unit #1 PSMC was reviewed. This test procedure will be signature table-top reviewed prior to commencement of tests.
Inputs for an update to the ACIS grounding diagram were developed during the ACIS Mid-Term review at MIT. Inputs to the grounding drawing will be incorporated during the next reporting period with an expected release soon after the first of the year.
Start of the PSMC EMI/EMC testing was delayed because of the Open Frame testing needed to validate procedures and to perform closed box performance measurements. It is anticipated that the testing will commence the first week in December.
The focus of Systems Engineering for the month of November was support for two meetings (Spacecraft Charging at TRW and ACIS Mid-Term Review at MIT), follow-up on the engineering estimate for Change Order #43, and a coordination of MIT comments to requirements and interface control documents. Additionally, there were final preparations for the pending EMI/EMC test of the PSMC.
Review by MIT of comments incorporation into the ACIS PTS Specification was completed this month. Redlines will be coordinated during the next reporting period with an anticipated release soon thereafter.
Update of the GSE CEI Specification is planned for completion next month with a draft out for review mid January 1996. Comments are due by late January with a final publication by early February 1996.
Revision D "Draft" of the Observatory to SI ICD (IF1-20) remains un-baselined. Preliminary cost and schedule impacts for the "Jiffy Pin" PIRN were developed. There have been several other Preliminary Interface Revision Notice (PIRN) changes identified as having impacts but without quantification of schedule or cost. In the aggregate, these changes have significant impact to the PTS hardware. These impacts will be gathered and provided when available.
Drafts of the ACIS PTS to DPS ICD and the ACIS FP to DH ICD have been reviewed by MIT. Written comments for incorporation are expected during the next reporting period. Final release of both ICDs has slipped to early January 1996.
During this reporting period, Systems Design activities supported Spacecraft Charging requirements being developed by TRW for the AXAF-I mission. A one day trip to attend a Spacecraft Charging TIM , a precursor to the EDI CDA at TRW, provided valuable information with respect to design requirements for ESD pulses on primary power and signal interfaces. The group continued support of the AXAF-I Weekly Action Item Tracking (WAIT) telecon on Wednesdays.
Developed additional task plan estimating data to assure validity of the "Stretch" engineering estimate.
Telecon discussions between MIT and LMA (Cambridge) personnel about test and ground system processing flows for the ACIS instrument hardware continued throughout November. LMA continues to review the ACIS component and system level test flows with respect to "Stretch" received from MIT. The various assembly flows are inter-related and require several more review and update cycles by MIT and LMA technical leads. Ongoing effort continues for requirements flowdown from the AXAF-IPD PRD to the ICD to ACIS CEI's and eventually to the LMA PTS Specification. This task is nearing completion with an anticipated presentation at the January 1996 MIT/LMA TIM.
Schedule remains the most significant management problem.
Electrical power requirements (Watts) are summarized in the following table:
|Peak power distribution
in Standby Mode
|Peak power distribution
in Max. Operating Mode
|Peak power distribution
in Bakeout Mode
|Peak power distribution in
Normal Operating Mode*
Note: Normal operating mode refers to the ACIS operating with six analog chains at full power, six front-end processors at full power, one back-end processor at full power and the focal plane temperature being maintained at -120 deg C.